OpenStudy (anonymous):

HELP ASAP!!!! PLEASE An exam consists of two parts, section x and section y. there can be a maximum of 95 questions. there must be at least 20 more questions in section Y than in section X. write a system of inequalities to model the number of questions in each of the two sections. then solve the system by graphing

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

well we know that \[x+y \le 95\] \[x+20 \le y\]\[x \ge 0\] and \[y \ge 0\] so I would have to go with b

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yep